I need help factoring x4 + x2y2 + y4 (please note that the 4's and 2's are exponents)
Answer by stanbon(75887)

help factoring x4 + x2y2 + y4
= x^4 + 2x^2y^2 + y^4 - x^2y^2
= (x^2+y^2)^2 - (xy)^2
= (x^2+y^2+xy)(x^2+y^2-xy)
===============
Cheers,
Stan H.